# Active Learning for Control-Oriented Identification of Nonlinear Systems

### v1

@@ 2 Problem Formulation
with state $X_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{X}}}$, input $U_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{U}}}$, and $d_{\mathsf{X}}$-dimensional noise $W_{t}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{w}^{2}I)$. For simplicity, we assume $X_{1}=0$. Here, $f$ is the dynamics function, which depends on an unknown parameter $\phi^{\star}\in\mathbb{R}^{d_{\phi}}$. We assume that there exists some positive $B$ such that $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$
@@
The superscript on the expectation denotes that the dynamics (1) are rolled out under parameter $\phi$, while the subscript denotes that the system is played in closed-loop under the feedback control policy $U_{t}=\pi_{t}(X_{1},U_{1},\dots,X_{t-1},U_{t-1},X_{t})$ for $t=1,\dots,T$. The learner follows a two step interaction protocol with an exploration phase, and an evaluation phase. In the exploration phase, the learner interacts with the system for a total of $N$ episodes, each consisting of $T$ timesteps, by playing exploration policies $\pi\in\Pi_{\exp}$. The policy class $\Pi_{\exp}$ is an exploration policy class, described in more detail below. The learner does not incur any cost during the exploration episodes, and seeks only to gain information about the system. After the $N$ interaction episodes, it uses the collected data to propose a policy $\hat{\pi}\in\Pi^{\star}$. The learner is then evaluated on the expected cost of the proposed policy on a new evaluation episode. In particular, it incurs cost $\mathcal{J}(\hat{\pi},\phi^{\star})$.
@@ 2.1 Certainty Equivalent Control
$\hat{\phi}\in\operatorname*{argmin}_{\phi\in\mathbb{R}^{d_{\phi}}}\sum_{n=1}^{N}\sum_{t=1}^{T}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\phi)\right\|^{2}.$ (3)
@@ 2.2 Assumptions
(Bounded Costs) There exists some $r_{\mathsf{cost}}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\mathsf{cost}}(\phi^{\star}))$, and all $\pi\in\Pi^{\star}$, we have $\mathbb{E}_{\pi,\phi}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]\leq L_{\mathsf{cost}}$.
Due to the smoothness assumptions on the dynamics (Assumption 2.1) and the policy class (Assumption 2.2), the above assumption holds if the stage costs $c_{t}$ are smooth.
We additionally suppose that $\Pi^{\star}$ is not overparametrized.
@@
The global minimum of $\mathcal{J}(\pi^{\theta},\phi^{\star})$, $\theta^{\star}(\phi^{\star})$, is unique and satisfies $\nabla^{2}_{\theta}\mathcal{J}(\pi^{\theta},\phi^{\star})|_{\theta=\theta^{\star}(\phi^{\star})}\succ 0$.
The above condition requires that there are a unique set of optimal control parameters for the true system. [^2] To gain an intuition for this condition, observe that it holds for linear quadratic regulators as long as the input penalty matrix is positive definite.
Under the above assumptions, we may show that the certainty equivalent policy is smooth near the optimal solution.
Lemma 2.1 (Proposition 6 of Wagenmaker et al. (2023)).
@@
While the Lojasiewicz assumption ensures that the data collected via the exploration policy $\pi^{0}$ is sufficient to identify the parameters, the rate of recovery may be slow. To bypass this limitation, we assume that some policy in the exploration class satisfies a persistence of excitation condition. Persistence of excitation for a nonlinear dynamical system involves the positive definiteness of the following matrix:
$\Sigma^{\pi}\triangleq\operatorname{\mathbb{E}}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right],$ (7)
where $Df(X_{t},U_{t},\phi^{\star})$ is the Jacobian of $f$ with respect to $\phi$ evaluated at $\phi^{\star}$.
@@
$\operatorname{\mathbb{E}}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right]\geq\lambda_{\min}^{\star}>0,$
Finally, we make an assumption that all models parameters are relevant to perfect control.
Assumption 2.7.
The model-task Hessian is positive definite: $\mathcal{H}(\phi^{\star})\succ 0$.
The above assumption is quite strong, and does not hold for common baseline problems such as the linear-quadratic regulator with fully unknown dynamics matrices. However, this assumption is made primarily for ease of exposition. It can be removed in the linear setting at the expense of an algorithm which is more complicated to express concisely (Wagenmaker et al., 2023). The same approach used to bypass this condition in the linear setting can be easily extended to the nonlinear setting considered in this paper. We do not formalize this here, although we will in future work.
@@ 3 Proposed Algorithm and Main Result
The above smoothness assumptions allow us to characterize the excess control cost of a policy synthesized via certainty equivalence applied to a parameter estimate $\hat{\phi}$, (5).
@@
 [^3] Thm. 1 of Wagenmaker et al. (2023) holds for systems with linear dependence on the unknown parameters; however it generalizes with minimal modification to the proof. See Appendix B.
Suppose Assumptions 2.1-2.4 hold. Let $r_{\theta}(\phi^{\star})$ be as defined in Lemma 2.1. Then for $\hat{\phi}\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
Footnote 3 informs us that the leading term of the excess cost is given by the parameter estimation error weighted by the model-task Hessian, $\left\|\hat{\phi}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}$.
@@
$\lim_{N\to\infty}\sqrt{N}\mathbf{FI}^{\pi}(\phi^{\star})^{1/2}(\hat{\phi}-\phi_{\star})\sim\mathcal{N}(0,I),$
where the Fisher information for the parameter $\phi$ under the policy $\pi$ for the model (1) is given as
$\mathbf{FI}^{\pi}(\phi)\!=\!\frac{\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}\!Df(X_{t},U_{t},\phi)^{\top}\!Df(X_{t},U_{t},\phi)\right]}{\sigma_{w}^{2}},$ (9)
and measures the signal-to-noise ratio of the data collected from an episode of interaction with the system under the exploration policy $\pi$.
@@
Suppose Assumption 2.1 holds. Let $H$ be a positive definite matrix and let $\delta\in(0,\frac{1}{4}]$. Consider the least squares estimate $\hat{\phi}$ determined from (3) using data collected from $N$ episodes via an exploration policy $\pi$ which is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz for some $\alpha\in(\frac{1}{4},\frac{1}{2}]$, and satisfies $\lambda_{\min}\left(\Sigma^{\pi}\right)>0$, with $\Sigma^{\pi}$ as defined in (7). There exists a polynomial $\texttt{poly}_{\alpha}$ depending only on $\alpha$ such that the following condition holds. With probability at least $1-\delta$,
$\left\|\hat{\phi}-\phi^{\star}\right\|_{H}^{2}$ $\leq 40\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(H\mathbf{FI}^{\pi}(\phi^{\star})^{-1})}{N}$ (10)
as long as
$N$ $\geq\mathsf{poly}_{\alpha}\bigg{(}T,L_{f},d_{\phi},d_{\mathsf{X}},\sigma_{w},\log N,\log\frac{1}{\delta},\log B_{f},C_{\mathsf{Loja}}\bigg{)}.$
By substituting the inequality (10) into the leading term of (8), one would expect that the the excess cost of deploying the certainty equivalent policy synthesized on a least squares estimate determined from data collected using policy $\pi$ is characterized by
@@
Our main result shows that the above intuition can be made rigorous through careful analysis and design of the exploration policy. It then proceeds to show that we can find an exploration policy approximately solving (12), even though the parameter $\phi^{\star}$ defining the exploration objective is unknown prior to experimentation.
To circumvent the issue of the unknown parameter $\phi^{\star}$, we consider a two step approach in which we first obtain a crude parameter estimate, and then refine it by playing a targeted exploration policy. Denote the crude estimate by $\hat{\phi}^{-}$. This parameter can be used to search for a policy that approximately solves the optimization problem in (12). The straightforward approach to do so is to solve the problem under the estimated parameter:
@@
While such a strategy may be effective if the coarse estimate is sufficiently close to the ground truth, it contends with the challenge that there is a distribution shift between the setting that the exploration policy is optimized for, and that which it is deployed in. In particular, $\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-1})$ pertains to the statistics from rolling out the dynamics under a parameter value different from the ground truth. We therefore instead consider the following problem:
$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(H(\hat{\phi}^{-})\Lambda_{\tilde{\pi}}^{-1}\right)\mbox{ where }$ (14)
$\Lambda_{\tilde{\pi}}\!=\!\operatorname{\mathbb{E}}_{\tilde{\pi}}^{\phi^{\star}}\!\left[\sum_{t=1}^{T}Df(X_{t},U_{t};\hat{\phi}^{-})^{\top}Df(X_{t},U_{t};\hat{\phi}^{-})\right]$ (15)
Note the difference between (13) and (14). In (13), the Fisher information is determined using an expectation under the the dynamics parameter estimate $\hat{\phi}^{-}$ (see (9)). In contrast, the expectation in (14) is evaluated under the true parameter $\phi^{\star}$. If (14) can be optimized over the exploration policies, it alleviates the aforementioned distribution shift issue.
A minimizer for the above objective (14) may be approximated via an online learning procedure, as was recently proposed by Wagenmaker et al. (2023) for models that are linear in the parameter. In particular, we leverage the dynamic optimal experiment design (DOED) procedure proposed in the aforementioned work. For the reader’s convenience, we state the algorithm and highlight several modifications in Section C.1. We refer to our modified version of the algorithm as $\texttt{DOED}_{+}$.
@@
The above discussion motivates Algorithm 1, named Active Learning for Control-Oriented Identification (ALCOI). The algorithm takes as input an initial policy satisfying the Lojasiewicz condition (Assumption 2.5), the exploration policy class, the target policy, the number of exploration rounds, and the desired confidence parameter $\delta$. Given these components, the algorithm proceeds in three stages. The first stage begins in 2 by playing the initial policy for one fourth of the exploration budget. In 3, it uses the collected data to derive a coarse estimate $\hat{\phi}^{-}$ for the unknown parameters by solving a least squares problem. Then, in 4 it constructs a confidence ball $\mathcal{B}^{-}$ centered at $\hat{\phi}^{-}$:
$\mathcal{B}^{-}$ =$=\bigg{\{}\phi:\left\|\phi-\hat{\phi}^{-}\right\|\leq\frac{R}{N^{\alpha}}\bigg{\}},$ (16)
where
$R=\left(\frac{2048\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(L_{f}TN\right)+\log\frac{1}{\delta}\right)\right)^{\alpha}.$ (17)
Next, the estimate $\hat{\phi}^{-}$ is used to construct the model-task Hessian at $\hat{\phi}^{-}$ as $\mathcal{H}(\hat{\phi}^{-})$. This is in turn used to define an exploration objective, $\Phi:\mathbb{R}^{d_{\phi}\times d_{\phi}}\to\mathbb{R}$ to be passed to the $\texttt{DOED}_{+}$ algorithm. This exploration objective is a regularized version of that in (14) which takes as input a Gram matrix $\Lambda$, and outputs $\operatorname{\mathrm{tr}}(\mathcal{H}(\hat{\phi}^{-})(\Lambda+\lambda I)^{-1})$. Here $\lambda>0$ is a regularizer chosen from the estimate of the model-task Hessian in 5. The quantity $B_{\Phi}$ is then defined from the model-task Hessian and the regularizer to characterize the smoothness of the objective $\Phi$. The exploration objective is passed to $\texttt{DOED}_{+}$ along with the coarse estimate $\hat{\phi}^{-}$, the ball $\mathcal{B}^{-}$ and the smoothness bound $B_{\Phi}$ in 8. This call to $\texttt{DOED}_{+}$ returns a collection of $\tilde{K}$ policies, where $\tilde{K}$ is a polynomial function of the supplied budget; see Section C.1 for further details. The mixture of the returned policies approximately minimizes the exploration objective $\Phi$ over Gram matrices of the form (15), thereby approximating a solution to (14). We therefore use them to design a mixture policy which plays any of the returned policies with equal probability, and plays the initial exploration policy $\pi^{0}$ with probability $\frac{1}{2}$. [^4] The initial policy $\pi^{0}$ included in $\pi_{\mathsf{mix}}$ to ensure that $\pi_{\mathsf{mix}}$ is a $(2C_{\mathsf{Loja}},\alpha)$-Lojasiewicz policy. This policy is run to collect data from the system, and obtain a fresh estimate $\hat{\phi}^{+}$ for $\phi^{\star}$. [^5] In practice, one would use all of the the data to estimate $\hat{\phi}^{+}$. However, using only the data from an independent rollout contributes only a constant factor to the rate, and simplifies the analysis. Finally, the estimate is used to synthesize the certainty equivalent policy as in (5).
Algorithm 1 $\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\delta)$
@@
1:Input: Initial policy $\pi^{0}$, exploration policy class $\Pi_{\exp}$, target policy class $\Pi^{\star}$, episodes $N$, confidence $\delta$.
@@
2: Play $\pi^{0}$ for $\lfloor\frac{N}{4}\rfloor$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},x_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{4}\rfloor}$.
@@
4: Set $\mathcal{B}^{-}$ according to (16).
5: Set $\lambda\leftarrow\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}$.
6: Define objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\hat{\phi}^{-})(\Lambda+\lambda I)^{-1})$.
7: Set objective smoothness bound $B_{\Phi}\leftarrow\frac{TL_{f}^{2}\left\|\mathcal{H}(\hat{\phi}^{-})\right\|}{\lambda^{2}}$.
8: Determine exploration policies as $\mathopen{}\left\{\pi^{k}\right\}\mathclose{}_{k=1}^{\tilde{K}}\leftarrow\texttt{DOED}_{+}(\Phi,\lfloor\frac{N}{4}\rfloor,\pi^{0},\hat{\phi}^{-},\mathcal{B}^{-},B_{\Phi})$.
9: Define $\pi_{\mathsf{mix}}$ which at the start of each episode plays $\pi^{0}$ with probability $\frac{1}{2}$, and for $i=1,\dots,\tilde{K}$ plays $\pi^{i}$ with probability $\frac{1}{2\tilde{K}}$.
10:Play $\pi_{\mathsf{mix}}$ for $\lfloor N/2\rfloor$ episodes, collecting data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{2}\rfloor}$.
11: Fit $\hat{\phi}^{+}$ by solving (3) with the data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{2}\rfloor}$.
12:Return: certainty equivalent policy $\hat{\pi}=\pi^{\star}(\hat{\phi}^{+})$
@@
Suppose $f$, $\pi^{0}$, $\Pi_{\exp}$, $\Pi^{\star}$ satisfy Assumptions 2.1-2.7. Let $\delta\in(0,\frac{1}{4}]$. Consider running Algorithm 1 to generate a control policy $\hat{\pi}$ as $\hat{\pi}=\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\delta)$. There exists a universal positive constant $c$ and polynomial function $\texttt{poly}_{\alpha}$ depending only on the Lojasiewicz parameter $\alpha$ such that the following holds true. With probability at least $1-\delta$ we have that
@@
$\mathcal{J}(\hat{\pi},\phi^{\star})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\leq c\frac{\log\frac{d_{\phi}}{\delta}}{N}\!\min_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\right),$
as long as
$N\!$ $\geq\!\mathsf{poly}_{\alpha}\bigg{(}\!T,L_{f},L_{\mathsf{cost}},L_{\pi^{\star}},d_{\phi},d_{\mathsf{X}},\frac{1}{\lambda_{\min}^{\star}},r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})$

### v2

@@ 2 Problem Formulation
with state $X_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{X}}}$, input $U_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{U}}}$, and $d_{\mathsf{X}}$-dimensional noise $W_{t}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{w}^{2}I)$ for some $\sigma_{w}>0$. For simplicity, we assume $X_{1}=0$. Here, $f$ is the dynamics function, which depends on an unknown parameter $\phi^{\star}\in\mathbb{R}^{d_{\phi}}$. We assume that there exists some positive $B$ such that $\left\|\phi^{\star}\right\|\leq B-1$ and $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$ satisfying $\left\|\phi\right\|\leq B$.
@@
The functions $c_{t}$ are known stage costs. The superscript on the expectation denotes that the dynamics (1) are rolled out under parameter $\phi$, while the subscript denotes that the system is played in closed-loop under the feedback control policy $U_{t}=\pi_{t}(X_{1},U_{1},\dots,X_{t-1},U_{t-1},X_{t})$ for $t=1,\dots,T$. The learner follows a two step interaction protocol with an exploration phase, and an evaluation phase. In the exploration phase, the learner interacts with the system for a total of $N$ episodes, each consisting of $T$ timesteps, by playing exploration policies $\pi\in\Pi_{\exp}$. The policy class $\Pi_{\exp}$ is an exploration policy class, described in more detail below. The learner does not incur any cost during the exploration episodes, and seeks only to gain information about the system. After the $N$ interaction episodes, it uses the collected data to propose a policy $\hat{\pi}\in\Pi^{\star}$. The learner is then evaluated on the expected cost of the proposed policy on a new evaluation episode. In particular, it incurs cost $\mathcal{J}(\hat{\pi},\phi^{\star})$.
@@ 2.1 Certainty Equivalent Control
$\hat{\phi}\in\operatorname*{argmin}_{\phi\in\mathbb{R}^{d_{\phi}},\left\|\phi\right\|\leq B}\sum_{n=1}^{N}\sum_{t=1}^{T}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\phi)\right\|^{2}.$ (3)
@@ 2.2 Assumptions
(Regular costs) The stage costs $c_{t}$ are three times differentiable and $\left\|D_{u}^{(i)}c_{t}(x,u)\right\|_{\mathsf{op}}\leq L_{\mathsf{cost}}$ for $i=1,\dots,3$, $\theta\in\mathbb{R}^{d_{\theta}}$, $(x,u)\in\mathbb{R}^{d_{\mathsf{X}}+d_{\mathsf{U}}}$. There exists some $r_{\mathsf{cost}}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\mathsf{cost}}(\phi^{\star}))$, and all $\pi\in\Pi^{\star}$, we have $\mathbb{E}_{\pi}^{\phi}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]\leq L_{\mathsf{cost}}$.
We additionally assume that the certainty equivalent policy is a smooth function of the dynamics parameter in a neighborhood around the optimal parameter.
@@
While the Lojasiewicz assumption ensures that the data collected via the exploration policy $\pi^{0}$ is sufficient to identify the parameters, the rate of recovery may be slow. To bypass this limitation, we assume that some policy in the exploration class satisfies a persistence of excitation condition. This condition can be expressed by first defining the Fisher information matrix for a parameter $\phi$ and a policy $\pi$ as
$\mathbf{FI}^{\pi}(\phi)\!\triangleq\!\frac{\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}\!Df(X_{t},U_{t},\phi)^{\top}\!Df(X_{t},U_{t},\phi)\right]}{\sigma_{w}^{2}},$ (7)
where $Df(X_{t},U_{t},\phi)$ is the Jacobian of $f$ with respect to $\phi$. The Fisher information measures the signal-to-noise ratio of the data collected from an episode of interaction with the system under exploration policy $\pi$. With this definition, persistance of excitation is equivalent to the positive definiteness of the matrix $\mathbf{FI}^{\pi}(\phi^{\star})$.
@@
$\mathbf{FI}^{\pi}(\phi^{\star})\succeq\mu I\succ 0.$
@@ 3 Proposed Algorithm and Main Result
The above smoothness assumptions allow us to characterize the excess control cost of a policy synthesized via certainty equivalence applied to a parameter estimate $\hat{\phi}$, (5). In particular, we extend a result from Wagenmaker et al. (2023) from the linear in parameters setting to the smooth nonlinear setting.
@@
Suppose Assumptions 2.1-2.4 hold. Let $r_{\theta}(\phi^{\star})$ be as defined in Assumption 2.4. Then for $\hat{\phi}\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
Lemma 3.1 informs us that the leading term of the excess cost is given by the parameter estimation error weighted by the model-task Hessian, $\left\|\hat{\phi}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}$.
@@
$\lim_{N\to\infty}\sqrt{N}\mathbf{FI}^{\pi}(\phi^{\star})^{1/2}(\hat{\phi}-\phi_{\star})\sim\mathcal{N}(0,I).$
@@
Suppose Assumption 2.1 holds. Consider the least squares estimate $\hat{\phi}$ determined from (3) using data collected from $N$ episodes via an exploration policy $\pi$ which is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz for some $\alpha\in(\frac{1}{4},\frac{1}{2}]$, and satisfies $\lambda_{\min}\left(\mathbf{FI}^{\pi}(\phi^{\star})\right)>0$, with $\mathbf{FI}^{\pi}(\phi^{\star})$ as defined in (7). Let $H$ be a positive definite matrix, $\beta$ a positive number satisyfing $\beta\leq\sigma_{w}^{2}\frac{\lambda_{\min}(\mathbf{FI}^{\pi}(\phi^{\star}))}{4},$ and $\delta\in(0,\frac{1}{4}]$. Then there exists a polynomial $\mathsf{poly}_{\alpha}$ depending on $\alpha$ such that the following condition holds. With probability at least $1-\delta$,
$\left\|\hat{\phi}-\phi^{\star}\right\|_{H}^{2}\leq 2(1+\xi)$ (9)
×$\times\left(\frac{\operatorname{\mathrm{tr}}(H\mathbf{FI}^{\pi}(\phi^{\star})^{-1})}{N}+2\frac{\left\|H\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|}{N}\log\frac{4}{\delta}\right),$
where $\xi=4\beta\left(\frac{1}{\sigma_{w}^{2}\lambda_{\min}(\mathbf{FI}^{\pi}(\phi^{\star}))}+d_{\phi}\right)$ as long as
$N$ $\geq\mathsf{poly}_{\alpha}\bigg{(}T,L_{f},\!d_{\phi},\!d_{\mathsf{X}},\!\sigma_{w},\!\log N,\!\log\frac{1}{\delta},\!\log\frac{B}{\sigma_{w}},\!C_{\mathsf{Loja}},\!\frac{1}{\beta}\bigg{)}.$
By substituting the inequality (LABEL:eq:_id_error) into the leading term of (8) and bounding $\left\|\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|\leq\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})$, one would expect that the the excess cost of deploying the certainty equivalent policy synthesized on a least squares estimate determined from data collected using policy $\pi$ is characterized by
@@
Our main result shows that the above intuition can be made rigorous through careful analysis and design of the exploration policy. It then proceeds to show that we can find an exploration policy approximately solving (11), even though the parameter $\phi^{\star}$ defining the exploration objective is unknown prior to experimentation. It is also necessary to address the fact that the model-task Hessian may not be positive definite. Thus optimizing the above objective could result in exploration policies which are not persistently exciting.
To circumvent the issue of the unknown parameter $\phi^{\star}$, we consider a two step approach in which we first obtain a crude parameter estimate, and then refine it by playing a targeted exploration policy. Denote the crude estimate by $\hat{\phi}^{-}$. This parameter can be used to search for a policy that approximately solves the optimization problem in (11). A straightforward approach to do so is to solve the problem under the estimated parameter:
@@
To address the issue of a model-task Hessian which is not positive definite, we introduce regularization into the exploration design. In particular, we set the exploration policy $\pi$ as
$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}((H(\hat{\phi}^{-})+\nu I)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}),$ (13)
for an appropriately chosen regularization parameter $\nu$.
@@
The above discussion motivates Algorithm 1, named Active Learning for Control-Oriented Identification (ALCOI). The algorithm takes as input an initial policy satisfying the Lojasiewicz condition (Assumption 2.5), the exploration policy class, the target policy, the number of exploration rounds, a parameter $\gamma\in(0,1)$ which controls the ratio of the exploration budget that the initial loja policy is played, the level of regularization $\nu$, and the precision of the optimization for the exploration policy $\varepsilon$. Given these components, the algorithm proceeds in three stages. The first stage begins in 2 by playing the initial policy for a portion of the exploration budget controlled by $\gamma$. In 3, it uses the collected data to derive a coarse estimate $\hat{\phi}^{-}$ for the unknown parameters by solving a least squares problem. Next, the estimate $\hat{\phi}^{-}$ is used to construct the model-task Hessian as $\hat{\phi}^{-}$ as $\mathcal{H}(\hat{\phi}^{-})$ and define an exploration objective $\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\lambda I\right)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}\right)$. Optimizing this objective to precision $\varepsilon$ over the class of exploration policies provides the policy $\pi_{\exp}$. This policy is run to collect data from the system, and obtain a fresh estimate $\hat{\phi}^{+}$ for $\phi^{\star}$. Finally, the estimate is used to synthesize the certainty equivalent policy as in (5).
Algorithm 1 $\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\gamma,\nu,\varepsilon)$
@@
1:Input: Initial policy $\pi^{0}$, exploration policy class $\Pi_{\exp}$, target policy class $\Pi^{\star}$, initial policy ratio $\gamma$, regularization parameter $\nu$, optimization precision $\varepsilon$.
@@
2: Play $\pi^{0}$ for $\lfloor\gamma N\rfloor$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\gamma N\rfloor}$.
@@
4: Determine exploration policy as
$\pi_{\exp}\in\Bigg{\{}\pi\in\Pi_{\exp}|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1}\right)$
$\leq(1+\varepsilon)\inf_{\tilde{\pi}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}\right)\Bigg{\}}.$
5: Define $\pi_{\mathsf{mix}}$ which at the start of each episode plays $\pi^{0}$ with probability $\gamma$, and $\pi_{\exp}$ with probability $1-\gamma$.
6: Play $\pi_{\mathsf{mix}}$ for $\lfloor(1-\gamma)N\rfloor$ episodes, collecting data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor(1-\gamma)N\rfloor}$.
7: Fit $\hat{\phi}^{+}$ by solving (3) with the data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor(1-\gamma)N\rfloor}$.
8:Return: certainty equivalent policy $\hat{\pi}=\pi^{\star}(\hat{\phi}^{+})$.
@@
Suppose $f$, $\pi^{0}$, $\Pi_{\exp}$, $\Pi^{\star}$ satisfy Assumptions 2.1-2.6. Let $\nu$ be a non-negative regularization parameter such that $\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu>0$. Let the optimization tolerance $\varepsilon\in(0,1/2)$ and the initial policy ratio $\gamma\in(0,1/2)$. Consider running Algorithm 1 to generate a control policy $\hat{\pi}$ as $\hat{\pi}=\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\gamma,\nu,\varepsilon)$.
@@
Let $\delta\in\big{(}0,\frac{1}{4}]$ be the failure probability, and $\beta\in\bigg{(}0,\frac{\mu\left(\lambda_{\min}(H(\phi^{\star}))+\nu\right)}{512d_{\phi}(\left\|H(\phi^{\star})\right\|+\nu)}\bigg{)}$ [^3] Recall that $\mu$ is the persistance of excitation parameter defined in Assumption 2.6. be a free parameter in the bound. There exists a polynomial function $\texttt{poly}_{\alpha}$ depending on the Lojasiewicz parameter $\alpha$ such that the following holds true. With probability at least $1-\delta$, it holds that
@@
$\mathcal{J}(\hat{\pi},\phi^{\star})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\!\leq\!(1\!+\!4\gamma)(1\!+\!\varepsilon)(1\!+\!\xi)\left(2+4\log\frac{4}{\delta}\right)$
×$\times\frac{\inf_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\right)}{N},$